OpenStudy (anonymous):

can anyone help me integrate (1+1/t)(1/t^2)

OpenStudy (anonymous):

I assume you are not allow to multiply out , right?

OpenStudy (anonymous):

I have to use substitution

OpenStudy (anonymous):

\[\left(1+\frac{1}{t}\right)\frac{1}{t^2}\]

OpenStudy (anonymous):

actually wait it is (1+1/t)^3(1/t^2)

OpenStudy (anonymous):

\[-t ^{-1}+(-t ^{-4}\div4)\]

OpenStudy (anonymous):

substitute u for 1/t

OpenStudy (anonymous):

simplify first, integrate separately.